# Fast 3D point clouds retrieval for Large-scale 3D Place Recognition

### v1

@@ 1 Introduction
Retrieval in 3D point clouds involves finding the most similar point clouds to a given query within a reference database of 3D point clouds. Research on retrieval within 3D point clouds is still in its early stages, with solutions mainly focused on the indexing phase (i.e. the production of point cloud descriptors) rather than the retrieval phase, which involves efficiently finding similar descriptors in the reference. The current state of the art confirms that the main application of retrieval in point clouds is geolocalization. Unlike approaches relying on visual place recognition, 3D LiDAR point clouds have the benefits of being invariant to lighting and weather conditions [13], making them particularly relevant for this purpose.
@@
The main objective of the article is to contribute to the retrieval phase, by proposing a scalable retrieval method, facing the size of the reference dataset and the dimensionality of the descriptors. We explore the application of Differentiable Search Index (DSI) [21] methods, originally developed for Information Retrieval in large text corpora, to the domain of 3D point clouds. DSI employs a single T5 (text-to-text transfer transformer) model to map string queries to relevant document identifiers (docids). To bridge the gap between the 3D point cloud queries and the textual input required for Information Retrieval, we explored existing Place Recognition methods, such as PointNetVLAD [22] and LoGG3D-Net [24], which have shown promising results in loop detection by encoding a point cloud into a global descriptor, a single feature vector that summarizes the entire point cloud geometry. Our model takes inspiration from Generative Image-to-text Transformer (GIT) [25] which is a Vision Transformer (ViT) used for image captioning. Fig. 1 illustrates the whole process, where the GIT image encoder is replaced with a point cloud encoder where the text decoder is trained to map 3D data to unique identifiers. The main contributions of this paper are:
@@ 2 Related work and main concepts
In this section, we present a state of the art on 3D point clouds retrieval in Section 2.1, followed by a closer look at the loss function used by LoGG3D-Net in Section 2.2. We present the concepts of Differentiable Search Index in 2.3 and how it can be adapted to 3D data with Vision-language model in Section 2.4.
@@ 2.1 Retrieval in 3D point clouds
Most of the approaches of retrieval in 3D point clouds encountered are applied to the problem of place recognition, i.e. the capability of determining the geolocation of a points cloud query from the most similar clouds retrieved in a geolocalized reference. First, the reference dataset is usually divided into multiple smaller samples, each of them being geolocalized and indexed with a descriptor. Second, the same type of descriptor is computed for the query and compared to those of the reference dataset. The best matches yield the most likely locations of the query.
@@
Global learned features-based. By contrast with handcrafted methods, learned point clouds descriptors have benefited from deep learning. A pioneering end-to-end learning-based method, PointNetVLAD [22], generates a global descriptor from point cloud data. It combines PointNet [16] for local feature extraction, with NetVLAD [1], a learnable CNN layer that aggregates the local features with into a single global descriptor. Since NetVLAD is a symmetric function, i.e, it is input order invariant, so can be applied to unordered point clouds.
@@
Multi-sourced global features. Cross-source 3D place recognition concerns the confrontation of point clouds acquired at different viewpoints, e.g. ground-based LiDAR scans with aerial ones. It presents significant challenges due to differences in coverage, density and noise patterns. To bridge the representation gaps of those types of acquisition, CrossLoc3D [7] introduces a cross-source 3D matching approach by learning a shared embedding space between aerial LiDAR database and ground LiDAR scans. It employs multi-scale sparse convolutions for feature selection, followed by a NetVLAD aggregator to obtain a unique descriptor. An iterative refinement step aligns the two embedding spaces into a shared one.
@@ 2.2 LoGG3D-Net
where $P_{q}$ is the query point cloud, $P_{p}$ is a positive one within $p=3m$ of the query, $P_{n}$ is a negative point cloud distant from $n=20m$ of the query, $P_{nbis}$ respects the negative condition with both $P_{p}$ and $P_{n}$. g($P_{*}$) is the final global descriptor vector, $N$ is the number of sampled negatives and $[]_{+}$ the hinge loss function, while $\alpha$ and $\beta$ are the loss margin of the quadruplet. For the retrieval step, the Cosine distance performs global descriptors retrieval by similarity between the query and the reference dataset items. In [24], previous entries adjacent to the query in time by less than $30s$ are excluded from the search to avoid matching with the same instance.
@@ 3.1 DSI-3D
Applying the DSI method requires converting 3D data into text form. This involves a two-step process: first, encoding the 3D point cloud into an embedding, and then passing this embedding through a multimodal transformer to generate captions. For the point cloud encoding, we choose LoGG3D-Net [24] due to its local consistency loss, which significantly enhances retrieval performances for place recognition, but other point cloud global descriptors could be used. The multimodal transformer chosen is GIT [25] for its streamlined architecture (see Section 2.4). By substituting the image encoder of GIT with a 3D encoder, we adapt it to generate captions that serve as docids for point clouds, derived from scene descriptors. For the training part, we use both the regular GIT loss and a variation of the language modeling $L_{m}$ loss (3) to be triplet and quadruplet losses $L_{tripl}$ and $L_{quad}$ (both in $\left[0,+\infty\right[$), as follows:
@@
With $L_{mpos}$, $L_{mneg}$ and $L_{mnegbis}$ representing the cross-entropy between the scene docid and output logits for a positive point cloud, a negative point cloud, and a second negative point cloud, respectively. $\alpha$ and $\beta$ are resp. fixed at 0.5 and 0.3. Finally, the retrieval step is performed with a beam search with a restricted vocabulary to meaningful indexes.
@@ 3.2 Docid representation
For the representation of the docid, we exclude Unstructured Atomic identifiers (see Section 2.3), as they are the least scalable option. For Naively Structured String identifiers, a natural analogy can be drawn with the naming of each point cloud in classical benchmarks with acquisition sequences (e.g. KITTI for mobile mapping), where each point cloud is named sequentially from 0 to $N$. The Semantically Structured identifiers strategy is also assessed. In this approach, neighboring point clouds with similar descriptors, and thus being in the same cluster, will share similar identifier prefixes; naturally, this method extends well to revisits.
@@
As an example, the docid returned for a point cloud $n$ with center coordinates $x_{n}$=’$1$’ and $y_{n}$=’$2$’ will be $docid$=’$12$’. This bijective mapping between point clouds and identifiers guarantees that each revisit of a scene will have a closely related identifier. We also propose a variant of this positional indexing, using space-filling curves, which map a multidimensional space onto a 1D curve, such as the Hilbert curve, known for its superior locality preservation [5]. After shifting and rescaling the $X_{N}$ and $Y_{N}$ coordinates into positive integers, they form a 2D grid where center coordinates of each point cloud are assigned a unique identifier derived from the Hilbert curve. For our experiments with the KITTI dataset, the Hilbert curve is applied with 17 iterations.
@@
In this section, we present the dataset, the evaluation criteria and the experimental setup chosen. Here, LoGG3D-Net serves both as our 3D encoder for DSI-3D as a baseline for comparison.
@@ 5.1 Loss function
Table 1 presents the $F1_{max}$ and $Hits@1$ scores obtained for the six original sequences on our model, with the GPS strategy, by exploiting the triplet and the quadruplet losses of Section 3.1 (again with a negative distance threshold of $3m$). We observe that the triplet loss achieves a marginally higher average $F1_{max}$ score, while the quadruplet one performs slightly better on the $Hits@1$ score. For the remaining experiments, we chose the quadruplet loss as it provides the best results on both metrics for the longest sequences 00 and 02 and with a shorter amount of training.
@@
F1max 00 02 05 06 07 08
$L_{tripl}$ 0.9775 0.9725 0.9946 0.9674 0.9954 0.9875
$L_{quad}$ 0.9809 0.9758 0.9891 0.9674 0.9955 0.9851
Hits@1 00 02 05 06 07 08
$L_{tripl}$ 0.9560 0.9465 0.9892 0.9369 0.9820 0.9755
$L_{quad}$ 0.9626 0.9507 0.9783 0.9369 0.9909 0.9705
@@ 5.2 Indexing strategies comparison
Table 2 and Table 3 resp. present the $F1_{max}$ scores and the $Hits@1$ scores obtained for the LoGG3D-Net baseline (without and with index LSH) and several variants of DSI-3D, namely: names ’Label’, ’Hierar’, ’GPS’ and ’Hilbert’ refer to the indexing strategies Naively Structured String identifiers, Semantically Structured identifiers, Positional Structured identifiers and Hilbert curve Positional Structured identifiers respectively. The sequence 22 was only trained with the Hilbert curve Positional Structured identifiers as it is the most promising indexing strategy.
@@
Secondly, among the various indexing strategies tested for DSI-3D, Hilbert curve indexing yields the best results, conducing to the conclusion that the strategies of the state of the art (’Label’ and ’Hierar’) can be outdated for retrieval dedicated to place recognition. This version does not achieve the same performance as the exact baseline LoGG3D-Net, which provides slightly better $Hits@1$ score, by an average improvement of 1.6% (median at 1.3%) ; for $F1_{max}$, ’Hilbert’ is less efficient of 0.8% on average (median at 0.5%). However, we observe the same behavior, as well as scores comparable to DSI-3D ones, for approximate retrieval when adding LSH256 to LoGG3D-Net. Globally, DSI-3D with Hilbert curve indexing achieves even a higher average $F1_{max}$ than LoGG3D-Net with LSH256 by 0.4% (median at 0.6%) and a better average $Hits@1$ by 0.1% (median at 0.6%).
As conclusion of this section, the behavior of DSI-3D, in terms of retrieval quality, can be considered at least as effective as a conventional state-of-the-art approximate approach. In the following section, we discuss its superiority facing the questions of complexity and retrieval time.
@@ 5.3 Complexity and retrieval time
In its current implementation [24], retrieval with LoGG3D-Net is performed over the entire reference dataset by sequential comparisons (first line of Tables 2 and 3), making retrieval time proportional to the dataset size ($n$) with a time complexity $O(n)$. The state-of-the-art solutions listed in Section 2.1 proceed in the same way. Similarity search in such high-dimensional spaces can be accelerated with indexes like the LSH [15], as experimented in the same Tables (lines 2 and 3), whose complexity largely differs from the variant used but is generally dominated by $O(n^{\rho})$ with $0\leq\rho\leq 1$, under the approximate retrieval assumption. In contrast, DSI-3D relies on model inference with constant $O(1)$ complexity, which is a significant advantage facing large datasets of point clouds. It also works under the approximate retrieval paradigm, but we have experimented in the previous section that the retrieval quality remains fully comparable with the ones of other solutions under the same paradigm.
@@
For DSI-3D indexed with Positional Structured identifiers, currently retrieval time is $0.45s$, by measuring the time it takes for a descriptor to pass through all GIT layers and return the docid. It is worth noting that the current code implementation used in the article is not yet optimized. Nevertheless, to give an insight on the current scalability of DSI-3D, we fitted a linear equation between the size of the reference datasets $N_{desc}$ and retrieval times $T$ with exact LoGG3D-Net ($T=0.67N_{desc}-0.68$), which allows to estimate that the inversion point occurs at $N_{desc}$ = 756.4e3 descriptors in favor of DSI-3D. According to the settings used, this amount roughly corresponds to a spatial coverage of 50km², at which point DSI-3D would become faster in its current not optimized implementation.
@@ 6 Conclusion
In this article, we have presented DSI-3D, a new indexing and retrieval approach for 3D point clouds, that drastically accelerates retrieval, compared to current literature which mainly focuses on point cloud description and performs retrieval following a linear time complexity as experimented here, it can be slightly improved with off-the-shelf index such as LSH. The proposal relies on the adaptation of the Derivative Search Index, but also introduces two new Positional Structured identifiers, fitted for the problem of place recognition. This article has laid the foundations of the approach, exhibiting performances similar to those of the state of the art in terms of retrieval quality. In the future, it will be necessary to continue the experiments, in particular by varying the model parameters further, and also by considering larger datasets to confirm its robustness in terms of the quality of the returned responses, its optimality in terms of response time, following a constant complexity, having been proven.

### v2

@@ 1 Introduction
Retrieval in 3D point clouds involves finding the most similar point clouds to a given query within a reference database of 3D point clouds. Research on retrieval within 3D point clouds is still in its early stages. Existing solutions mainly focus on the indexing phase (i.e. the production of point cloud descriptors) rather than the retrieval phase, which involves efficiently finding similar descriptors in the reference. The current state of the art confirms that the main application of retrieval in point clouds is geolocalization. Unlike approaches relying on visual place recognition, 3D LiDAR point clouds have the benefits of being invariant to lighting and weather conditions [13], making them particularly relevant for this purpose.
@@
The main objective of the article is to contribute to the retrieval phase, by proposing a scalable retrieval method, facing the size of the reference dataset and the dimensionality of the descriptors. We explore the application of Differentiable Search Index (DSI) [21] methods, originally developed for Information Retrieval in large text corpora, to the domain of 3D point clouds. DSI employs a single T5 (text-to-text transfer transformer) model to map string queries to relevant document identifiers (docids). To bridge the gap between the 3D point cloud queries and the textual input required for Information Retrieval, we explored existing Place Recognition methods, such as PointNetVLAD [22] and LoGG3D-Net [24]. These methods have shown promising results in loop detection by encoding a point cloud into a global descriptor, a single feature vector that summarizes the entire point cloud geometry. Our model takes inspiration from Generative Image-to-text Transformer (GIT) [25] which is a Vision Transformer (ViT) used for image captioning. Fig. 1 illustrates the whole process, where the GIT image encoder is replaced with a point cloud encoder where the text decoder is trained to map 3D data to unique identifiers. The main contributions of this paper are:
@@ 2 Related work and main concepts
In this section, we present a state of the art on 3D point clouds retrieval in Section 2.1, followed by a closer look at the loss function used by LoGG3D-Net in Section 2.2. We present the concepts of Differentiable Search Index in 2.3 and how it can be adapted to 3D data with a Vision-language model in Section 2.4.
@@ 2.1 Retrieval in 3D point clouds
Most of the approaches of retrieval in 3D point clouds encountered are applied to the problem of place recognition. This involves determining the geolocation of a points cloud query from the most similar clouds retrieved in a geolocalized reference. First, the reference dataset is usually divided into multiple smaller samples, each of them being geolocalized and indexed with a descriptor. Second, the same type of descriptor is computed for the query and compared to those of the reference dataset. The best matches yield the most likely locations of the query.
@@
Global learned features-based. By contrast with handcrafted methods, learned point clouds descriptors have benefited from deep learning. A pioneering end-to-end learning-based method, PointNetVLAD [22], generates a global descriptor from point cloud data. It combines PointNet [16] for local feature extraction, with NetVLAD [1], a learnable CNN layer that aggregates local features into a single global descriptor. Since NetVLAD is a symmetric function, i.e, it is input order invariant, so can be applied to unordered point clouds.
@@
Multi-sourced global features. Cross-source 3D place recognition concerns the confrontation of point clouds acquired at different viewpoints, e.g. ground-based LiDAR scans with aerial ones. It presents significant challenges due to differences in coverage, density and noise patterns. To bridge the representation gaps of those types of acquisition, CrossLoc3D [7] introduces a cross-source 3D matching approach. It learns a shared embedding space between aerial LiDAR database and ground LiDAR scans. It employs multi-scale sparse convolutions for feature selection, followed by a NetVLAD aggregator to obtain a unique descriptor. An iterative refinement step aligns the two embedding spaces into a shared one.
@@ 2.2 LoGG3D-Net
where $P_{q}$ is the query point cloud, $P_{p}$ is a positive one within $p=3m$ of the query, $P_{n}$ is a negative point cloud distant from $n=20m$ of the query, $P_{nbis}$ respects the negative condition with both $P_{p}$ and $P_{n}$. g($P_{*}$) is the final global descriptor vector. $N$ is the number of sampled negatives and $[]_{+}$ the hinge loss function, while $\alpha$ and $\beta$ are the loss margin of the quadruplet. For the retrieval step, the Cosine distance performs global descriptors retrieval by similarity between the query and the reference dataset items. In [24], previous entries adjacent to the query in time by less than $30s$ are excluded from the search to avoid matching with the same instance.
@@ 3.1 DSI-3D
Applying the DSI method requires converting 3D data into text form. This involves a two-step process: first, encoding the 3D point cloud into an embedding, and then passing this embedding through a multimodal transformer to generate captions. For the point cloud encoding, we choose LoGG3D-Net [24] due to its local consistency loss, which significantly enhances retrieval performances for place recognition, but other point cloud global descriptors could be used. The multimodal transformer chosen is GIT [25] for its streamlined architecture (see Section 2.4). By substituting the image encoder of GIT with a 3D encoder, we adapt it to generate captions that serve as docids for point clouds, derived from scene descriptors. In order to reinforce the training by increasing the similarity between positive pairs as in LoGG3D-Net, we use both the regular GIT loss and a variation of the language modeling $L_{m}$ loss (3) to be triplet and quadruplet losses $L_{tripl}$ and $L_{quad}$ (both in $\left[0,+\infty\right[$). These losses are computed using tuples of point clouds: $(P_{q},P_{p},P_{n})$ for the triplet one, and $(P_{q},P_{p},P_{n},P_{nbis})$ for the quadruplet one, as follows:
@@
With $L_{mpos}$, $L_{mneg}$ and $L_{mnegbis}$ representing the cross-entropy loss between the docid of a query point cloud $P_{q}$ and output logits for a positive point cloud $P_{p}$, a negative point cloud $P_{n}$, and a second negative point cloud $P_{nbis}$, respectively. $P_{nbis}$ forms a negative pair with both the query $P_{q}$ and the first negative $P_{n}$ ; $\alpha$ and $\beta$ are resp. fixed at 0.5 and 0.3. Finally, the retrieval step is performed with a beam search with a restricted vocabulary to meaningful indexes.
@@ 3.2 Docid representation
For the representation of the docid, we exclude Unstructured Atomic identifiers (see Section 2.3), as they are the least scalable option. For Naively Structured String identifiers, a natural analogy can be drawn with the naming of each point cloud in classical benchmarks with acquisition sequences (e.g. KITTI for mobile mapping), where each point cloud is named sequentially from 0 to $N$. The Semantically Structured identifiers strategy is also assessed. In this approach, neighboring point clouds with similar descriptors, and thus being in the same cluster, will share similar identifier prefixes. Naturally, this method extends well to revisits.
@@
As an example, the docid returned for a point cloud $n$ with center coordinates $x_{n}$=’$1$’ and $y_{n}$=’$2$’ will be $docid$=’$12$’. This bijective mapping between point clouds and identifiers guarantees that each revisit of a scene will have a closely related identifier. We also propose a variant of this positional indexing, using space-filling curves, which map a multidimensional grid space onto a 1D curve, such as the Hilbert curve, known for its superior locality preservation [5]. After shifting and rescaling the $X_{N}$ and $Y_{N}$ center coordinates into positive integers, each point cloud center is assigned a unique identifier derived from the Hilbert curve. This identifier corresponds to the distance the curve must travel to reach point $(X_{N},Y_{N})$, ensuring that the spatially close points have similar index representations. Empirically, for our experiments with the KITTI dataset, we apply the Hilbert curve with 17 iterations, resulting in a grid resolution of $2^{17}\times 2^{17}$.
@@ 4 Experimental Setup
In this section, we present the dataset, the evaluation criteria and the experimental setup chosen. Here, LoGG3D-Net serves both as our 3D encoder for DSI-3D and as a baseline for comparison.
@@ 5.1 Loss function
Table 1 presents the $F1_{max}$ and $Hits@1$ scores obtained for the six original sequences on our model, with the GPS strategy, by exploiting the triplet and the quadruplet losses of Section 3.1 (again with a negative distance threshold of $3m$). We observe that the triplet loss achieves a marginally higher average $F1_{max}$ score, while the quadruplet one performs slightly better on the $Hits@1$ score. For the remaining experiments, we chose the quadruplet loss, as it provides the best results on both metrics for the longest sequences 00 and 02.
@@
F1max 00 02 05 06 07 08 mean
$L_{tripl}$ 0.9775 0.9725 0.9946 0.9674 0.9954 0.9875 0.9825
$L_{quad}$ 0.9809 0.9758 0.9891 0.9674 0.9955 0.9851 0.9823
Hits@1 00 02 05 06 07 08 mean
$L_{tripl}$ 0.9560 0.9465 0.9892 0.9369 0.9820 0.9755 0.9644
$L_{quad}$ 0.9626 0.9507 0.9783 0.9369 0.9909 0.9705 0.9650
@@ 5.2 Indexing strategies comparison
Table 2 and Table 3 resp. present the $F1_{max}$ scores and the $Hits@1$ scores obtained for the LoGG3D-Net baseline (without and with index LSH) and several variants of DSI-3D. The names ’Label’, ’Hierar’, ’GPS’ and ’Hilbert’ refer to the indexing strategies Naively Structured String identifiers, Semantically Structured identifiers, Positional Structured identifiers and Hilbert curve Positional Structured identifiers respectively. Sequence 22 was only trained with the Hilbert curve Positional Structured identifiers as it is the most promising indexing strategy.
@@
Secondly, among the various indexing strategies tested for DSI-3D, Hilbert curve indexing yields the best results. This leads to the conclusion that the strategies of the state of the art (’Label’ and ’Hierar’) can be outdated for retrieval dedicated to place recognition. This version does not achieve the same performance as the exact baseline LoGG3D-Net, which provides a slightly better $Hits@1$ score, by an average improvement of 1.6% (median at 1.3%) ; for $F1_{max}$, ’Hilbert’ is less efficient of 0.8% on average (median at 0.5%). However, we observe the same behavior, as well as scores comparable to DSI-3D ones, for approximate retrieval when adding LSH256 to LoGG3D-Net. Globally, DSI-3D with Hilbert curve indexing achieves even a higher average $F1_{max}$ than LoGG3D-Net with LSH256 by 0.4% (median at 0.6%) and a better average $Hits@1$ by 0.1% (median at 0.6%).
In conclusion of this section, the behavior of DSI-3D, in terms of retrieval quality, can be considered at least as effective as a conventional state-of-the-art approximate approach. In the following section, we discuss its superiority facing the questions of complexity and retrieval time.
@@ 5.3 Complexity and retrieval time
In its current implementation [24], retrieval with LoGG3D-Net is performed over the entire reference dataset by sequential comparisons (first line of Tables 2 and 3). This makes the retrieval time proportional to the size of the dataset ($n$) with a time complexity $O(n)$. The state-of-the-art solutions listed in Section 2.1 proceed in the same way. The similarity search in such high-dimensional spaces can be accelerated with indexes such as the LSH [15], as experimented in the same Tables (lines 2 and 3). The complexity largely differs from the variant used but is generally dominated by $O(n^{\rho})$ with $0\leq\rho\leq 1$, under the approximate retrieval assumption. In contrast, DSI-3D relies on model inference with constant $O(1)$ complexity, which is a significant advantage facing large datasets of point clouds. It also works under the approximate retrieval paradigm. However, as we have experimented in the previous section, the retrieval quality remains fully comparable to the ones of other solutions under the same paradigm.
@@
For DSI-3D indexed with Positional Structured identifiers, currently retrieval time is $0.09s$, by measuring the time it takes for a descriptor to pass through all GIT layers and return the docid. It is worth noting that the current code implementation used for the article is not yet optimized. Nevertheless, to give an insight on the current scalability of DSI-3D, we fitted a linear equation between the size of the reference datasets $N_{desc}$ and retrieval times $T$ with exact LoGG3D-Net ($T=0.67N_{desc}-0.68$). This allows us to estimate that the inversion point occurs at $N_{desc}$ = 135.3e3 descriptors in favor of DSI-3D. According to the settings used, this amount roughly corresponds to a spatial coverage of 10km², at which point DSI-3D would become faster in its current not optimized implementation.
@@ 6 Conclusion
In this article, we have presented DSI-3D, a new indexing and retrieval approach for 3D point clouds. It drastically accelerates retrieval, compared to current literature which mainly focuses on point cloud description and performs retrieval following a linear time complexity as experimented here, it can be slightly improved with off-the-shelf index such as LSH. The proposal relies on the adaptation of the Differentiable Search Index, but also introduces two new Positional Structured identifiers, fitted for the problem of place recognition. This article has laid the foundations of the approach, exhibiting performances similar to those of the state of the art in terms of retrieval quality. In the future, it will be necessary to continue the experiments, in particular by varying the model parameters further, and also by considering larger datasets to confirm its robustness in terms of the quality of the returned responses, its optimality in terms of response time, following a constant complexity, having been proven.